OpenStudy (anonymous):

The sum of -15/32 and 7/32 is _____.

OpenStudy (alexandervonhumboldt2):

this is easy: -15/32 + 7/32=(-15+7)/32=\[\frac{ -15+7 }{ 32 }\]

OpenStudy (alexandervonhumboldt2):

simplify this now

OpenStudy (anonymous):

1/4